At a physician's office, patients sit in the Waiting Room until it is their turn (first-come, first-served) to be seen by the doctor. There is only one doctor on duty, and the doctor can help only one patient at a time. Based on her long-run rate of client service, this doctor typically spends 30 minutes per patient. Whenever the doctor finishes her session with one patient, she immediately begins a new session with the next patient.

When you arrive in the Waiting Room, the doctor has just begun her session with a patient, and you see that there are 4 other patients already waiting.

In this situation, it is reasonable to model your time (hours) in the Waiting Room as an Erlang random variable with:

λ=?

w=?
Answer #1

From the given information, patients sit in the waiting room until it is their turn on First- come First serve basis. There iWait in the Queue is calculated as Then the waiting in the system is calculated as W =W +- - = 0.5 + 1 / 2 = 0.5+0.5 = 1 Ther